What is the frequency of a photon with a wavelength of 1040 nm. Report your answers to three significant digits.


1
Expert's answer
2021-07-28T02:49:02-0400

frequency, "f=c/\lambda=(3\times 10^8)/(1.04\times 10^{-6})=2.89\times 10^{14}Hz"